# KiVA: Kid-inspired Visual Analogies for Testing Large Multimodal Models

### v2

@@ 2 Related Work
Another specific class of benchmarks tests generalization and reasoning within abstract puzzle grids. These include Abstraction and Reasoning Corpus (ARC) (Chollet, 2019; Moskvichev et al., 2023; Mitchell et al., 2023); a direct translation of RPMs-based human evaluation has previously been applied to models by (Ahrabian et al., 2024) and (Huang et al., 2024) (also see prior benchmarks (b) and (c) in Figure 2). However, the stimuli are simple, monotonic shapes like squares and circles, lacking real-world complexity and variability. Moreover, they primarily focus on convoluted pattern recognition and logical sequencing without incorporating real-world contexts, thereby neglecting more fundamental visual cognition domains that even human children are capable of.
In summary, although numerous benchmarks assess advanced visual capabilities in large multimodal models, none evaluate abilities that are demonstrably present in young children, such as drawing analogies about transformations of simple real objects.
@@ 4.1 Results
Model performance depends on the visual domain and correlates with human performance. Overall, models are better at classifying and describing color and size transformations than transformations in other domains (Figures 4 and 5), which involve more discrete and local processing than the other domains (Zeki et al., 1991; Zeng et al., 2020). In KiVA-adults, the best-performing model GPT-o1 nears adult performance only in the color domain (Figure 5). Models are less able to specify what changed within the visual domains of rotation, reflection, and number and consequently also did not perform well in extrapolations for those domains. In contrast, children and adults generally show similar performance across visual domains, with children performing slightly worse on rotation compared to other domains. Adults’ response times correlate with GPT-o1’s error scores (1-Accuracy) across the domains in visual extrapolation of KiVA, as demonstrated in Figure 7. What is cognitively demanding to humans is also more computationally challenging for GPT-o1.
@@
Verbal questions facilitate visual extrapolation in humans but the effects are less clear in GPT-o1. We included verbal questions to reveal the step in the reasoning process where models might fail when making visual analogies. To assess the effects of verbal questions on subsequent visual extrapolation, we tested another 200 adults, 20 children and the best-performing model, GPT-o1, on a visual-extrapolation-only version of KiVA, removing the verbal questions to replicate previous visual analogy benchmarks. We focused on testing the three more challenging visual domains of KiVA, number, rotation and reflection.Without verbal questions, adults demonstrated similar accuracy but significantly slower response times, whereas children performed worse in extrapolation. The effects are less clear for GPT-o1: it performed equally well in the number domain, better at extrapolating object rotations but worse at extrapolating reflections when asked to reason about what changed and how it changed beforehand (Figure 9). While our verbal questions facilitate humans’ visual extrapolation performance, it is possible that reasoning models like GPT-o1 already reason about “what changed” and “how it changed” independently of our verbal queries. Future work should further explore the effects of guiding questions and chain-of-thought on reasoning models. At the same time, it may also be possible to solve KiVA without language, as in the case of a Large Vision Model (Bai et al., 2024) that is trained in the complete absence of linguistic data (see Section B.4).
@@ Acknowledgments
We are grateful to Joe Heyward, Viorica Patraucean, Mariel Goddu and Jefferson Ortega and the participants of the AI, Psychology, and Neuroscience workshop at the Simons Institute for discussion, to participants and their parents, local early childhood centers, and UC Berkeley undergrads who assisted in human data collection: Alexis Davis, Janna Umagat, Kate Choi, Kaydee Manikhong, Linda Marie Trevino, Nitya Sriram, Nora Chen, Ray Huang and Shivalika Jhabua. This project was supported by Meta-BAIR Commons, CIFAR Catalyst Award: Causally guided exploration in children & AI, and ONR MURI Self-Learning Perception Through Real-World Interaction.
@@ B.2 Models’ extrapolation performance based on previous verbal reasoning
Furthermore, we report models’ extrapolation performance conditional on succeeding (green) or failing (red) at the previous steps of verbal reasoning in Figure 12. GGPT-o1 exhibits significantly higher visual extrapolation accuracy when its preceding verbal reasoning is correct across all transformation domains, whereas GPT-4V shows this benefit only in the color and size domains. In other words, successful visual extrapolation is contingent on solving verbal classification or specification correctly when models are solving KiVA above chance level. Meanwhile, there is no conditional dependence of prior verbal reasoning on subsequent visual extrapolation in LLaVA-1.5 and MANTIS, and they also perform no better than chance level on KiVA

### v3

@@ 2 Related Work
Another specific class of benchmarks tests generalization and reasoning within abstract puzzle grids. These include the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and ConceptARC (Moskvichev et al., 2023; Mitchell et al., 2023); a direct translation of RPMs-based human evaluation has previously been applied to models by (Ahrabian et al., 2024) and (Huang et al., 2024) (see these prior benchmarks in Figure 2). However, the stimuli are simple, monotonic shapes like squares and circles, lacking real-world complexity and variability. Moreover, they emphasize complex pattern recognition and logical sequencing without real-world context—neglecting basic visual cognition skills even children possess—and this limited scope may render them unsuitable for training data that typically covers a much broader range of real-world visuals.
In summary, although many benchmarks assess advanced visual capabilities in large multimodal models, none evaluate visual cognition that is clearly exhibited by young children—such as predicting simple transformations of real-world objects—or use children as a baseline for comparison.
@@ 4.1 Results
Model performance depends on the visual domain and correlates with human performance. Overall, models are better at classifying and describing color and size transformations than transformations in other domains (Figures 4 and 5), which involve more discrete and local processing than the other domains (Zeki et al., 1991; Zeng et al., 2020). In KiVA-adults, the best-performing model GPT-o1 nears adult performance only in the color domain (Figure 5). Models are less able to specify what changed within the visual domains of rotation, reflection, and number and consequently also did not perform well in extrapolations for those domains. In contrast, children and adults generally show similar performance across visual domains, with children performing slightly worse on rotation compared to other domains. Children’s error scores (1-Accuracy) and adults’ response times correlate with GPT-o1’s error scores in the visual extrapolation of KiVA, as demonstrated in Figure 6. What is cognitively demanding to humans is also more computationally challenging for GPT-o1.
@@
Verbal questions facilitate visual extrapolation in humans but the effects are less clear in GPT-o1. We included verbal questions to reveal the step in the reasoning process where models might fail when making visual analogies. To assess the effects of verbal questions on subsequent visual extrapolation, we tested another 200 adults, 20 children and the best-performing model, GPT-o1, on a visual-extrapolation-only version of KiVA, removing the verbal questions to replicate previous visual analogy benchmarks. We focused on testing the three more challenging visual domains of KiVA, number, rotation and reflection.Without verbal questions, adults demonstrated similar accuracy but significantly slower response times, whereas children performed worse in extrapolation. The effects are less clear for GPT-o1: it performed equally well in the number domain, it is better at extrapolating object rotations but worse at extrapolating reflections when asked to reason about what changed and how it changed beforehand (Figure 9). While our verbal questions facilitate humans’ visual extrapolation performance, it is possible that reasoning models like GPT-o1 already reason about “what changed” and “how it changed” independently of our verbal queries. Future work should further explore the effects of guiding questions and chain-of-thought on reasoning models. At the same time, it may also be possible to solve KiVA without language, as in the case of a Large Vision Model (Bai et al., 2024) that is trained in the complete absence of linguistic data (see Section B.4).
@@ Acknowledgments
We are grateful to Joe Heyward, Viorica Patraucean, Mariel Goddu, Jefferson Ortega and the participants of the AI, Psychology, and Neuroscience workshop at the Simons Institute for discussion, to participants and their parents, local early childhood centers, and UC Berkeley undergrads who assisted in human data collection: Alexis Davis, Janna Umagat, Kate Choi, Kaydee Manikhong, Linda Marie Trevino, Nitya Sriram, Nora Chen, Ray Huang, Shivalika Jhabua and Weiyin Gao. This project was supported by Meta-BAIR Commons, CIFAR Catalyst Award: Causally guided exploration in children & AI, and ONR MURI Self-Learning Perception Through Real-World Interaction.
@@ B.2 Models’ extrapolation performance based on previous verbal reasoning
Furthermore, we report models’ extrapolation performance conditional on succeeding (green) or failing (red) at the previous steps of verbal reasoning in Figure 12. GPT-o1 exhibits significantly higher visual extrapolation accuracy when its preceding verbal reasoning is correct across all transformation domains, whereas GPT-4V shows this benefit only in the color and size domains. In other words, successful visual extrapolation is contingent on solving verbal classification or specification correctly when models are solving KiVA above chance level. Meanwhile, there is no conditional dependence of prior verbal reasoning on subsequent visual extrapolation in LLaVA-1.5 and MANTIS, and they also perform no better than chance level on KiVA.